### v3

# Exploratory Notes on Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

Abstract
I present a coordinate-free, symbolic framework for deciding whether a given set of polygonal faces can form a closed, genus-zero polyhedral surface and for predicting how such a surface could be decomposed into internal tetrahedra. The method uses only discrete incidence variables, such as the number of internal tetrahedra $T$, internal gluing triangles $N_{i}$, and internal triangulation segments $S_{i}$, and applies combinatorial feasibility checks before any geometric embedding is attempted.
For polyhedra in normal form, I record exact incidence identities linking $V,E,F$ to a flatness parameter $S:=\sum_{f}(\operatorname{deg}f-3)$, and I identify parity-sensitive effects in $E$, $F$, and $S$. The external identities and parity-sensitive bounds hold universally for genus-0 polyhedral graphs. For internal quantities, I prove exact relations $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (with $S_{i}$ taken to be the number of interior edges) and obtain restricted linear ranges within a shell–aligned ladder subclass (SALT), where at most one interior edge is introduced per layer. Consequently, I propose a symbolic workflow that yields rapid pre-checks for structural impossibility, reducing the need for costly geometric validation in computational geometry, graphics, and automated modeling.
Tagging of Results:
  • Proposition / Lemma: Proved in this paper or from established literature in genus-zero polyhedra in normal form.
  • Conjectural Proposition: Believed true, supported by examples, but no general proof.
  • Heuristic Proposition: Empirical or pattern-based rule, not formally proved.
  • Observation: Fact specific to worked examples, not claimed to hold generally.
Author’s Note: This document represents original, exploratory research conducted independently by the author. It does not reflect the views or positions of any affiliated institution. The author does not claim to be a professional mathematician and welcomes constructive feedback, critical review, and potential collaboration to refine or expand upon the ideas presented.
1 Introduction and Background
This paper introduces a purely combinatorial framework for assessing the feasibility of polyhedral enclosures in three-dimensional space, based solely on symbolic logic and integer arithmetic. Unlike traditional geometric methods that rely on coordinates, angles, or embeddings, this approach evaluates the structural validity of polygonal configurations through discrete operations such as triangulation segment counts, vertex valency rules, and symbolic angle units.
Scope and fixed assumptions: Throughout this work, I limit attention to genus-zero (simply connected) surfaces, with no geometric embeddability tests (i.e., no attempt to prove realizability in $\mathbb{R}^{3}$), and without consideration of higher-genus or non-manifold topologies. Unless stated otherwise, all internal “range” results are given under the SALT+MIE restriction (Definition 2.1), i.e. normal form with at most one new interior segment per layer, introduced only when necessary. All identities and bounds presented are symbolic rather than metric, and some are conjectural or empirically validated rather than proved in full generality.
This work is exploratory and intended as a conceptual and symbolic approach to studying polyhedral closure. The methodology emerged from empirical analysis of polyhedral forms and was later formalized with computational assistance. It seeks to answer a longstanding question: when can a set of polygons form a valid 3D enclosure, and how can such enclosures be systematically analyzed or decomposed without geometric tools?
The approach is related in spirit to Ziegler’s symbolic treatment of polytopes [10] and to combinatorial efforts in tetrahedral decomposition (e.g., [4]), but differs in its exclusive reliance on integer logic and symbolic topological consistency. A classical result by Steinitz [8] shows that a graph is the 1-skeleton of a convex 3-polyhedron if and only if it is planar and 3-connected, providing the foundational combinatorial criterion for convex polyhedral graphs. While many known polyhedral validation tools rely on coordinate-based embeddings or optimization algorithms, the framework presented here derives topological invariants and structural bounds without any reference to spatial realization.
Recent extensions of Steinitz’s theorem have been explored for ball polyhedra [1], and Gallier and Quaintance [5] provide a contemporary combinatorial-topological survey covering shellings, Euler–Poincaré relationships, and triangulations. From a computational perspective, Below, De Loera, and Richter-Gebert [3] proved that finding a triangulation with the minimum number of tetrahedra is NP-complete. In contrast, the present method uses empirical observations to guide symbolic enumeration and narrow down the space of combinatorial configurations that may support enclosure and decomposition.
New contributions: A classical genus-zero identity (recorded here with a short derivation) for polyhedra in normal form is
$E=3V-6-S,\quad F=2V-4-S,$
where the flatness parameter
$S=\sum_{f}(\operatorname{deg}f-3)$
measures the total deviation from pure triangulation.
Combining these with the constraint that all vertices have degree at least 33 yields sharp bounds:
Even $V$:
$0\leq S\leq\frac{3V}{2}-6,\quad\frac{3V}{2}\leq E\leq 3V-6,\quad\frac{V}{2}+2\leq F\leq 2V-4.$
Odd $V$:
$0\leq S\leq\frac{3V}{2}-\tfrac{13}{2}$
$\frac{3V+1}{2}\leq E\leq 3V-6,\quad\frac{V}{2}+\tfrac{5}{2}\leq F\leq 2V-4.$
All integer $S$ values within these ranges are conjecturally realizable (Heuristic Proposition 3.2). We give explicit families realizing many values (e.g., stacked triangulations followed by controlled coplanarity merges), and exhaustively confirm small $V$ cases, but I do not claim a full classification here. – For even $V$, this includes purely cubic graphs or pentagonal–trivalent graphs (e.g., cube, dodecahedron, tetrahedron). – For odd $V$, $S_{\max}$ occurs when all but one vertex are trivalent and the remaining vertex is 4-valent.
In all cases, $S$ subtracts exactly from a “triangulation budget” of 3V63V-6 edges.
A parity-sensitive behavior is observed: while Euler’s characteristic $\chi=V-E+F$ is invariant under $V$-parity changes, the structures of $E$, $F$, and $S$ differ in predictable ways between even and odd vertex counts.
Building on this, I propose an extended Euler-type heuristic:
$V-E+F=2(T-N_{i}+S_{i}),$
linking external incidence data to internal decomposition variables: $T$ = internal tetrahedra, $N_{i}$ = internal gluing triangles, $S_{i}$ = internal triangulation segments.
This identity reduces to the classical Euler formula when $T-N_{i}+S_{i}=1$ (as in minimal or marching tetrahedralizations). Although not yet proved as a general necessity theorem, it functions as a strong symbolic consistency check.
Finally, empirical enumeration of symbolic decompositions reveals additional internal constraints on $(T,N_{i},S_{i})$, as well as distribution-dependent effects such as “lettucing,” in which evenly distributed flatness changes achievable internal configurations compared to concentrated flatness.
This document is structured as both a conceptual foundation and a practical toolbox for symbolic polyhedral reasoning. The core objective is to determine when a collection of polygonal faces, without any geometric embedding, can enclose a valid 3D volume, and if so, what internal structure it might support.
To ensure clarity, Table 1 summarizes the key variables and symbolic terms used throughout this document. These variables describe either the external surface of a polyhedron or structures required for its internal decomposition into tetrahedra.
Symbol Meaning
Surface / External Structure Variables
$V$ Number of vertices
$E$ Number of edges
$F$ Number of faces
$S$ External non-intersecting triangulation segments (used to triangulate surface); also used throughout as a discrete flatness measure
NN Total internal angle units on the external surface (1 unit $\pi$ radians $180^{\circ}$)
$F_{\triangle}$ Number of boundary triangular faces resulting from the triangulation of the entire polyhedral surface
$M$ External angle fullness units (derived from polygon types)
$a,b,c,\ldots$ Face types: triangle (3 sides), quadrilateral (4), pentagon (5), etc.
$a^{\prime},b^{\prime},c^{\prime},\ldots$ Vertex types: valency 3 (3 incident edges), valency 4, valency 5, etc.
Internal Decomposition Variables
$T$ Number of internal tetrahedra (used to fill the volume)
$N_{i}$ Number of internal gluing angle units (each unit == one triangle)
$S_{i}$ Number of interior edges (internal non-intersecting triangulation segments). In normal form there are no interior vertices, hence $S_{i}\equiv E_{i}$
$E_{i}$ (Alias) Number of interior edges; in normal form $E_{i}=S_{i}$
$V_{i}$ Internal vertices
Table 1: Key variables used in the symbolic framework. Surface/external variables describe the boundary structure, while internal variables capture the combinatorial scaffolding of tetrahedral decompositions. This grouping highlights the two structural layers that the framework relates through symbolic identities.
The variables above are not derived from physical measurements or coordinates, but from symbolic relationships among discrete structural components. For example, angle units are treated symbolically as multiples of $\pi$, and surface structures are quantified via segment and gluing unit counts rather than metric angles or lengths.
The auxiliary variables NN, $M$, $S$, $N_{i}$, and $S_{i}$ do not appear in the classical Euler–Steinitz framework. They are introduced here to support symbolic decomposition bookkeeping and to connect surface composition to volumetric tetrahedralization. N represents the total internal angle units on the external surface, with one unit defined as $\pi$ radians ($180^{\circ}$ ). $M$ is derived from polygon types and measures “external angle fullness” in the same units. $S$ counts the non-intersecting triangulation segments added to polygonal faces to produce a fully triangulated surface and also serves as a discrete flatness measure. $N_{i}$ counts the number of internal gluing angle units, each corresponding to a shared triangular face between two tetrahedra. $S_{i}$ counts the internal non-intersecting triangulation segments required to complete a volumetric mesh during decomposition. None of these auxiliary quantities are edges in the polyhedron’s graph; they exist only in the symbolic model and may not correspond to realizable geometric features.
1.1 Illustrative Examples of $N_{i}$, $S$, and $S_{i}$
Rectangular Pyramid ($S=1$):
A five-vertex polyhedron with a quadrilateral base and one apex point. To triangulate the external surface, a single diagonal segment divides the base into two triangles, yielding $S=1$.
Figure 1: Rectangular pyramid illustrating the flatness measure $S$. The single internal diagonal on the quadrilateral base (orange) is the unique surface triangulation segment needed to obtain a fully triangulated boundary, hence $S=1$. This corresponds to one coplanarity merge relative to the maximally triangulated case and, by Proposition 3.3, decreases both $F$ and $E$ by 11 while preserving $V$ and $\chi=2$.
Triangular Bipyramid ($N_{i}=1$):
Composed of two tetrahedra joined along a shared triangular face, this structure has one gluing triangle (blue): $N_{i}=1$.
Figure 2: Triangular bipyramid obtained by gluing two tetrahedra along a common triangular face (highlighted in blue). This shared face corresponds to exactly one internal gluing triangle, hence $N_{i}=1$ in the symbolic model. It provides the simplest example of how $N_{i}$ is counted: each such internal face removes two boundary triangles in the fully triangulated representation, linking interior adjacency directly to external face counts via Proposition 5.4.
Irregular Octahedron ($S_{i}=E_{i}=1$, $N_{i}=4$):
When regular, constructed by gluing two rectangular pyramids base-to-base. In the case of this irregular octahedron, the internal meshwork between the glued tetrahedra must be constructed by adding one internal segment functioning as an internal edge, yielding $S_{i}=E_{i}=1$. Once included, four gluing triangles would emerge, resulting in $N_{i}=4$.
Figure 3: Irregular octahedron constructed by gluing two quadrilateral-based pyramids along their bases. The orange segment is an internal triangulation segment ($S_{i}=1$) required to fully mesh the internal volume into tetrahedra without adding new vertices. This configuration produces four internal gluing triangles ($N_{i}=4$), shown in blue and yellow, illustrating how $S_{i}$ and $N_{i}$ together describe the combinatorial scaffolding of the internal decomposition.
2 Methods and Heuristics
This section introduces both the proved combinatorial identities and the empirical heuristics that underlie the framework, along with the triangulation logic that directs it towards constructing a polyhedron’s internal mesh.
2.1 Logical separation of results
Two classes of structural constraints are used in this work:
  1. 1.
    Proved genus-zero identities: These are exact, rigorously derived equalities and bounds that hold for all genus-zero polyhedra in normal form. Chief among them is the flatness identity
    $E=3V-6-S,\quad F=2V-4-S,$
    where $S=\sum_{f}(\operatorname{deg}f-3)$ measures the total deviation from triangulation. Interpreting $S=\sum_{f}(\operatorname{deg}f-3)$ as a flatness deficit, the identities imply the sharp ranges
    $\begin{cases}\dfrac{3V}{2},&V\ \text{even},\\[4.0pt] \dfrac{3V+1}{2},&V\ \text{odd}\end{cases}\;\;\leq\;E\;\;\leq\;3V-6,$
    $\begin{cases}\dfrac{V}{2}+2,&V\ \text{even},\\[4.0pt] \dfrac{V}{2}+\dfrac{5}{2},&V\ \text{odd}\end{cases}\;\;\leq\;F\;\;\leq\;2V-4,$
    $0\;\;\leq\;S\;\;\leq\;\begin{cases}\dfrac{3V}{2}-6,&V\ \text{even},\\[4.0pt] \dfrac{3V}{2}-\tfrac{13}{2},&V\ \text{odd}.\end{cases}$
    (Here the lower bound on $E$ uses $\operatorname{deg}(v)\geq 3$.) For each admissible $S$, $E$ and $F$ decrease by one from their triangulated values.
  2. 2.
    Empirical and conjectural constraints: These are patterns and formulas suggested by extensive manual and computational enumeration but not proved in general. A key example is the extended Euler-type identity
    $V-E+F=2(T-N_{i}+S_{i}),$
    where $T$ is the number of internal tetrahedra, $N_{i}$ is the number of internal gluing triangles, and $S_{i}$ is the number of internal triangulation segments. This identity is proved for certain construction sequences but remains a conjectural proposition in full generality.
In what follows, the derivation of heuristic bounds and decomposition strategies is always annotated with the appropriate tag (Heuristic, Proposition, or Conjectural proposition).
2.2 Empirical observation and validation
Through iterative experimentation, I tracked the effects of adding or removing one structural element at a time. On the surface, I examined vertices, edges, faces, and triangulation segments; internally, I observed how tetrahedra filled the polyhedron and how their gluing faces were defined by internal triangulation segments. I recorded all countable variables and analyzed their relationships. These empirical observations guided both the decomposition strategy and the derivation of formulas describing how both external and internal structures scale with vertex count and triangulation complexity (i.e., surface flatness). Extremal values of these variables were described as functions of $V$, and sequences were defined to characterize how they change relative to one another with increasing flatness for fixed $V$.
2.3 Surface flatness estimation
To quantify the flatness of a polyhedron’s surface, I introduce a discrete measure $S$, defined as the total number of triangulation segments added to convert all non-triangular external faces into triangles. Flatness in this context refers to the extent to which vertices on a given face are coplanar. The fewer triangular faces a polyhedron has initially, the more triangulation is required, and the higher its total flatness measure $S$.
Each polygonal face is triangulated by drawing non-intersecting segments between the nearest connectable vertices. The cumulative count of these added segments on all external faces constitutes the flatness measure $S$. When all faces of a polyhedron are triangles, $S=0$ and $F\equiv F_{\max}$. This discrete flatness measure is now linked directly to the proved genus-zero identity above, serving both as a predictor of external structure and as a measure of internal decomposition complexity.
2.4 Internal mesh construction from surface triangulation
Unlike the external surface of a polyhedron, which is both observable and abstractable with consistency, the internal structure must be inferred. Consequently, the number of tetrahedra required to tetrahedralize a polyhedron, such as a cube, depends critically on the configuration of face cuts. For example, as noted by Timalsina and Knepley [9], a cube may be divided into 5, 6, or 12 tetrahedra depending on the chosen strategy: five tetrahedra in an ideal face-aligned configuration, six using the marching tetrahedra algorithm, and twelve when a Steiner point is introduced at the cube’s center.
Since the goal of this exploration is to maximally preserve the original topology of the polyhedron, I adopt a triangulated, face-aligned tetrahedralization strategy. For this purpose, the internal mesh is constructed solely from external constraints, namely, the original vertices and triangulated faces.
To support this construction, internal triangulation segments are introduced to define the internal gluing triangles along which tetrahedra are joined. Each internal segment contributes once to the structural count $S_{i}$, and each internal triangle used for gluing contributes once to the structural count $N_{i}$, regardless of how many tetrahedra share them. Together, these inferred internal edges and triangle faces form a combinatorial scaffold that enables complete internal tetrahedralization while preserving the external surface geometry.
2.5 Normal-form restrictions on interior structure
Restricting attention to normal form constructions: all vertices lie on the boundary of the enclosing polygonal face, and no two edges intersect except at shared endpoints. In particular, interior vertices $V_{i}$ (vertices not incident to the boundary) cannot arise under these constraints, so $V_{i}\equiv 0$ in all cases considered. Moreover, every interior edge $E_{i}$ is necessarily an internal triangulation segment $S_{i}$, each belonging to exactly two distinct triangular faces. These conditions ensure that the triangulation is a planar straight-line graph in the classical sense, with all internal segments disjoint except at endpoints.
2.6 Tetrahedral decomposition and trivalent vertex prioritization
Original edges, surface triangulation segments, internal gluing triangles, and internal triangulation segments constitute the structure necessary to decompose the polyhedron into tetrahedra. These tetrahedra are joined along fully triangulated internal faces and, unless fully embedded, are partially or fully coned outward toward external triangulated faces. The coning of fully embedded tetrahedra is progressively revealed as the surrounding shell of tetrahedra is “chipped away” one at a time.
Tetrahedra are incrementally removed by eliminating vertices. Priority is given to trivalent vertices, those connected to exactly three edges, since each such vertex typically defines a single tetrahedron when coned to its surrounding triangular link. This targeted removal minimizes combinatorial complexity, which can occur when higher-valency vertices are eliminated without the right internal triangulation, potentially leading to more complex subdivisions and increased tetrahedral counts. When it is necessary to remove more complex vertices or parts of the polyhedron, the separation must occur across a single plane dividing coplanar vertices. Failing to respect such planar separations can result in the bisection of edges and faces, thereby increasing internal complexity.
This approach differs from geometric decompositions which rely on slicing through faces or bisecting edges to create tetrahedral meshes. In contrast, the method presented here preserves polygonal integrity: despite triangulation, all faces remain whole, with fixed vertex sets; no edge bisection or midpoint insertion is permitted. Internal tetrahedra are constructed exclusively through explicit gluing at existing vertices, maintaining a combinatorially pure structure.
Definition 2.1 (Shell–aligned ladder tetrahedralizations (SALT)).
In essence, I restrict attention to genus–0, normal–form decompositions built by peeling a single outer shell of tetrahedra at a time (a layer). Each layer is created by coning existing boundary triangles and, only if necessary, introducing at most one new interior segment between boundary vertices to complete gluing across that layer. No interior vertices (Steiner points) and no edge bisections are allowed. Internal adjacencies occur only along triangular faces.
Definition 2.2 (Minimal Interior Edge (MIE) rule).
A new interior segment is introduced only when strictly necessary to complete the current layer; otherwise none is added.
This SALT+MIE restriction excludes dense interior diagonals between boundary vertices (e.g. pulling triangulations of cyclic polytopes), which can create $\Theta(V^{2})$ tetrahedra without adding vertices. Our class supports the “ladder/shell” behavior analyzed in this paper.
2.7 Validations of observed relations by Euler’s characteristic
Euler’s characteristic is used here to validate the internal consistency of the framework.
The first validation is purely surface-based and corresponds to Proposition 3.1 together with Corollary 3.2. Beginning from the classical Euler characteristic and, in the genus-zero case, using incidence counting, I derive the proved flatness identity
$E=3V-6-S,\qquad F=2V-4-S,$
together with its associated extremal bounds. This establishes the exact relationship between $V$, $E$, $F$, and $S$ for genus-zero polyhedra in normal form.
The second validation links surface quantities to internal decomposition variables. Using Proposition 5.4, which expresses the boundary counts in terms of $T$ (internal tetrahedra), $N_{i}$ (internal gluing triangles), and $S_{i}$ (internal triangulation segments), I obtain the extended Euler-type heuristic
$V-E+F=2(T-N_{i}+S_{i}).$
This reduces to the classical value $\chi=2$ because $T-N_{i}+S_{i}=1$ along that ladder. Outside those sequences the identity is treated as a conjectural proposition pending a general proof.
2.8 Results taxonomy
Throughout the paper, statements are tagged as Heuristic, Proposition, or Conjectural proposition. Heuristic marks empirically validated bounds or patterns supported by computations and worked examples but not proved in full generality. Proposition denotes identities for which a complete proof is provided under stated hypotheses. Conjectural proposition designates candidate general identities whose proof would require showing that certain construction sequences cover all admissible decompositions.
3 Vertex-Driven External Polyhedral Properties: Intermediate Flatness Realizability and Parity Effects
I establish extremal bounds on external combinatorial quantities ($F$, $E$, and $S$) directly as functions of the vertex count $V$. These bounds emerge from Euler’s characteristic combined with flatness-based adjustments, and serve as the foundation for linking external structure to the internal decomposition results in later sections. This section proves that every integer flatness value in the admissible range $0\leq S\leq S_{\max}$ is realized by some genus-zero polyhedron in normal form. The realizability result is stated formally in Proposition 3.2.
The derivation proceeds from a given number of vertices $V\in\mathbb{N}$, showing how faces, edges, and triangulation segments co-evolve under Euler-consistent constraints. I examine parity effects in the incidence structure. Although Euler’s characteristic $\chi=V-E+F$ is invariant under vertex-parity changes, quantities such as $E$, $F$, and $S$ can vary systematically between even and odd $V$. These parity-sensitive patterns have implications for both surface enumeration and internal tetrahedral decomposition strategies. From this, I obtain a feasible range of structural properties for a closed 3D polyhedron and assess their validity using Euler’s formula.
3.1 Vertex Basis and Initial Angle Unit Calculation
Let $V\geq 4$ be the number of vertices in the polyhedron. I define:
  • NN: the total number of angle units on the external surface, where one unit = 180°, i.e., the internal angle of a triangle.
  • I assume full triangulation of the surface, where polygonal faces are split into triangles using non-intersecting internal segments through alternating vertices.
Then the total number of surface angle units is:
N=2(V2).N=2(V-2).
Before stating the general formulas, I note two standard planar-graph facts for the maximally triangulated case ($S=0$):
  1. 1.
    Each face is a triangle, so $3F_{\max}=2E_{\max}$. Combining this with Euler’s formula $V-E_{\max}+F_{\max}=2$ gives:
    $F_{\max}=2(V-2),\quad E_{\max}=3(V-2).$
  2. 2.
    A coplanarity merge of two adjacent triangles into a polygonal face removes exactly one face and one edge, leaving $V$ unchanged. If $S$ counts the number of such merges from the maximally triangulated case, then:
    $F=F_{\max}-S,\quad E=E_{\max}-S.$
Proposition 3.1 (External structure formulas).
For any simple, 3-connected planar graph on a genus-0 surface, with no digonal faces and in normal form with $V$ vertices, $E$ edges, $F$ faces, and flatness measure $S$ (the number of coplanarity merges from the maximally triangulated surface), the following identities hold:
$E=3V-6-S,\qquad F=2V-4-S.$
Proof.
For a genus-0 polyhedron in normal form, consider first the triangulated case ($S=0$). In this case every face has degree 33, so the incidence relation is
$3F_{\max}=2E_{\max}.$
Combining this with Euler’s formula $V-E_{\max}+F_{\max}=2$ yields
$F_{\max}=2V-4,\qquad E_{\max}=3V-6.$
Now, increasing the flatness parameter $S$ by one corresponds to replacing a triangular face by a face of degree 44; this removes exactly one edge and one face from the triangulation. More generally, each unit of $S$ reduces both $E$ and $F$ by exactly one compared to the triangulated case. Therefore
$F=F_{\max}-S=2V-4-S,$
$E=E_{\max}-S=3V-6-S,$
as claimed. ∎
Lemma 3.1 (Lower bound edges from trivalence).
Let $G$ be the 11-skeleton of a genus–0 polyhedron (simple, planar, and 33-connected) with $V\geq 4$ vertices and $E$ edges. Then
$E\;\geq\;\left\lceil\frac{3V}{2}\right\rceil,$
with equality when:
  1. 1.
    $V$ is even and all vertices are trivalent (degree 33); or
  2. 2.
    $V$ is odd and all but one vertex are trivalent, the remaining vertex being 44-valent.
Proof.
Every vertex has degree at least 33, so by the handshake lemma
$2E=\sum_{v\in V}\operatorname{deg}(v)\geq 3V,$
which yields $E\geq\lceil 3V/2\rceil$.
If $V$ is even and all vertices have degree 33, then $2E=3V$ and $E=3V/2$, attaining the bound.
If $V$ is odd and all but one vertex are trivalent and the remaining vertex is 44-valent, then 2E=3(V1)+4=3V+12E=3(V-1)+4=3V+1, so $E=(3V+1)/2$, again attaining the bound.
In the symbolic flatness framework, both equality cases correspond to $S=S_{\max}$ and $F=F_{\min}$ (see Corollary 3.2). ∎
Corollary 3.1 (Parity restriction on the minimal edge count).
Let $G$ be the 11–skeleton of a genus–0 polyhedron in normal form (simple, planar, 33–connected) with $V$ vertices and $E$ edges. If $V$ is odd, then $E\neq\frac{3V}{2}$ by integrality. Hence
$E_{\min}=\left\lceil\frac{3V}{2}\right\rceil=\frac{3V+1}{2}.$
Remark.
For odd $V$, the degree multiset achieving $E_{\min}$ necessarily has one 44–valent vertex and all others trivalent (by the handshake lemma). Realization for all odd $V$ is a separate existence question; examples exist for many $V$ (e.g., small explicit polyhedral graphs), but a full classification is beyond our scope.
Corollary 3.2 (Relating maximal flatness and minimal faces).
For $V\geq 4$, let $E_{\min}=\lceil 3V/2\rceil$. Then with the formulas from Proposition 3.1:
$S_{\max}=3V-6-E_{\min}=\begin{cases}\frac{3V}{2}-6,&\text{$V$ even},\\[2.0pt] \frac{3V}{2}-\tfrac{13}{2},&\text{$V$ odd},\end{cases}\qquad F_{\min}=2-V+E_{\min}=\begin{cases}\frac{V}{2}+2,&\text{$V$ even},\\[2.0pt] \frac{V}{2}+\tfrac{5}{2},&\text{$V$ odd}.\end{cases}$
Moreover, as $S$ increases by 11 (flattening step), both $E$ and $F$ decrease by 11, and for $k=0,1,\dots,S_{\max}$ one has E=3V6k,F=2V4k.E=3V-6-k,\;F=2V-4-k.
Remark.
On parity and the “half” constants. The halves (e.g., $S_{\max}=\tfrac{3V}{2}-\tfrac{13}{2}$ for odd $V$) come from the integer lower bound $E_{\min}=\lceil 3V/2\rceil$. Writing
$S_{\max}=3V-6-E_{\min}$
gives $S_{\max}=\tfrac{3V}{2}-6$ when $V$ is even and $S_{\max}=\tfrac{3V}{2}-\tfrac{13}{2}$ when $V$ is odd; these evaluate to integers at the respective parities. The corresponding formulas for $F_{\min}=2-V+E_{\min}$ have the same source.
Proof.
Substitute $E_{\min}$ into the formulas in Proposition 3.1. ∎
Heuristic Proposition 3.2 (Flatness values in the admissible range can be realized).
Fix $V\geq 4$. For every integer $S$ with $0\leq S\leq\frac{3V}{2}-6$ when $V$ is even and $0\leq S\leq\frac{3V}{2}-\tfrac{13}{2}$ when $V$ is odd, one can construct a genus-0 polyhedron in normal form (equivalently, a simple 33–connected planar graph) with
$E=3V-6-S,\qquad F=2V-4-S.$
In particular, every integer flatness value in the admissible range is realized.
Proof sketch.
Start from any maximally planar (triangulated) 33–connected planar graph on $V$ vertices (e.g., a stacked triangulation), which has $S=0$, E=3V6E=3V-6, F=2V4F=2V-4. A “coplanarity merge” of two adjacent boundary triangles corresponds combinatorially to deleting their shared edge, which decreases both $E$ and $F$ by 11 and keeps $V$ fixed. In a stacked triangulation one can choose a set of V4V-4 edges (e.g., along successive stacked faces) whose deletion preserves simplicity, planarity, and 33–connectivity; performing $S$ such merges yields a graph with the claimed $(E,F)$. ∎
3.2 External Angle Unit Bounds
From the flatness-based identities in Proposition 3.1, and holding $V$ fixed, the external angle unit range can be expressed as:
$M_{\max}=10(V-2),\qquad M_{\min}=\begin{cases}4V+6,&\text{$V$ odd},\\ 4V+4,&\text{$V$ even}.\end{cases}$
3.3 Configuration Series over Flatness
From Proposition 3.1, each unit decrease in the flatness measure $S$ increases both $F$ and $E$ by 11. Starting from the extremal values in Corollary 3.2, I have:
$S_{k}=S_{\max}-k,\quad F_{k}=F_{\min}+k,\quad E_{k}=E_{\min}+k,$
for $k=0,1,\dots,S_{\max}$.
3.4 Counting Distinct Combinatorial Configurations
By combining Proposition 3.1 and Corollary 3.2, the total number of distinct $(F,E)$ configurations attainable by varying $S$ over its admissible range is:
$\text{Combinations}=S_{\max}+1=\begin{cases}\frac{3}{2}V-\tfrac{11}{2},&\text{$V$ odd},\\ \frac{3}{2}V-5,&\text{$V$ even}.\end{cases}$
Heuristic 3.1 (External bounds and parity effects).
The expressions for the extremal ranges in §3 (including $F_{\min}$, $E_{\min}$, $S_{\max}$ and the parity-sensitive offsets) follow from Proposition 3.1 together with Lemma 3.1 and Corollary 3.1. They are supported by exhaustive checks for small $V$ and by constructive families achieving the bounds, but are not proved to be tight in full generality.
Angle units.
One “angle unit” equals the interior angle of a triangle, i.e., $180^{\circ}$. On a fully triangulated boundary, each triangle contributes exactly one unit and
$N=F_{\triangle}=F+S.$
Under an interior decomposition into $T$ tetrahedra with $N_{i}$ internal gluing triangles, counting triangle units gives
$N=4T-2N_{i},$
since each tetrahedron contributes four triangle units and every internal gluing triangle removes two boundary triangle units by pairing. These two expressions for NN are consistent with Proposition 5.4 and Proposition 3.1: indeed $F+S=4T-2N_{i}$. This connection between external angle counts and internal tetrahedral composition will be used in §5 to link internal volumetric parameters directly to external face and edge counts.
3.5 Triangulation Sequences and the Invariance of Euler’s Characteristic
Proposition 3.3 (Triangulated extrema and invariance of $\chi$).
For a genus-zero polyhedral surface composed entirely of triangles,
$F_{\max}=2(V-2),\qquad E_{\max}=3(V-2).$
If $S$ denotes the number of coplanarity merges of adjacent triangles (each merge decreases both $F$ and $E$ by 11 while $V$ is fixed), then for all $S$ in the admissible range,
$F(S)=F_{\max}-S,\qquad E(S)=E_{\max}-S,\qquad V-E(S)+F(S)=2.$
Proof.
Each triangle has three edges and each edge is shared by two triangles, so $2E_{\max}=3F_{\max}$. Substituting $E_{\max}=\tfrac{3}{2}F_{\max}$ into Euler’s formula $V-E_{\max}+F_{\max}=2$ yields $F_{\max}=2(V-2)$ and $E_{\max}=3(V-2)$. A coplanarity merge of two adjacent triangles removes one surface edge and one triangle, so $F\mapsto F-1$ and $E\mapsto E-1$ with $V$ unchanged. Hence VE+FV-E+F is invariant and equal to 22 for all admissible $S$. ∎
This result shows that the maximal triangulated case ($S=0$) fixes $F$ and $E$ uniquely in terms of $V$, and that the Euler characteristic is preserved under a sequence of coplanarity merges parameterized by $S$. In our framework, $S$ also functions as a discrete measure of flatness and appears in parity-sensitive bounds in this section, and will reappear in linking external and internal structures (§5). Appendix A lists worked examples for $V=4$ to $V=7$ illustrating these sequences and confirming VE+F=2V-E+F=2 in all cases.
4 Internal Decomposition and Tetrahedral Estimation
Section 4 shifts to internal volumetric decomposition and presents a clean identity linking tetrahedra, gluing planes, and internal segments. It first explores the internal volumetric structure of valid polyhedra, defined by a given vertex count $V$ and the surface properties derived in previous sections. Specifically, it aims to estimate the number or range of internal tetrahedra ($T$) that can fill a polyhedron, based on its triangulated faces and edge configurations. This continues the logic of the preceding analysis, shifting from surface validation to interior decomposition. The empirical formulas presented here result from the strict application of the tetrahedral decomposition process outlined in the Methods section.
4.1 Internal Structure Variables
Let a polyhedron with $V\geq 4$ vertices be internally composed of:
  • $T$: Number of tetrahedra,
  • $S_{i}$: Number of internal non-intersecting triangulation segments to fully triangulate internal gluing surfaces,
  • $N_{i}$: Number of internal gluing triangle units.
4.2 Lower bounds and restricted upper bounds (SALT+MIE)
Under SALT with the Minimal Interior Edge rule (at most one new interior edge per layer), the exact identities of Propositions 5.25.3 imply
$T_{\min}=V-3\quad(S_{i}=0),\qquad T=V-3+S_{i},\qquad N_{i}=2T-V+2=V-4+2S_{i}.$
Because at most one interior edge is introduced per layer and the first and last layers cannot host a new interior edge, one has $0\leq S_{i}\leq V-5$, whence
$T_{\max}=V-3+(V-5)=2(V-4),\qquad S_{i,\max}=V-5,\qquad N_{i,\max}=V-4+2(V-5)=3(V-4)-2.$
4.3 SALT ladder family
Within SALT+MIE, a transparent subfamily is the “ladder” construction where successive nontriangular boundary faces are filled one by one. For $k=0,\ldots,V-5$,
$T_{k}=V-3+k,\quad N_{i,k}=V-4+2k,\quad S_{i,k}=k,$
and $T_{k}-N_{i,k}+S_{i,k}=1$ holds exactly.
4.4 Count of Valid Internal Configurations
$\text{Count of valid }(T,N_{i},S_{i})\text{combinations}=V-4$
Heuristic 4.1 (Internal ranges and configuration ladder).
The ranges in §4 for $(T_{\min},T_{\max})$, $(N_{i,\min},N_{i,\max})$, and $S_{i,\max}$, together with the ladder
$T_{k}=V-3+k,\quad N_{i,k}=V-4+2k,\quad S_{i,k}=k\quad(0\leq k\leq V-5),$
match all worked examples and constructions tested to date, but are presently justified empirically rather than by a complete classification.
4.5 External Constraints on Internal Structure
While for polyhedra of minimal vertex count, surface flatness $S$ typically correlates with increased constraint on internal decomposition, restricting the range of valid $(T,N_{i},S_{i})$ configurations, this relationship becomes unreliable at higher vertex counts. In essence, internal complexity is not solely determined by how flat a polyhedron’s surface is, but also by how that flatness is distributed.
For $S=0$, the full tetrahedral range is always accessible, as no surface constraint exists.
At higher values of $S$, internal structure is expected to compress. However, I observe that when flatness is **distributed across many external faces**, the volume cannot be enclosed efficiently with shallow tetrahedral configurations alone. This leads to the appearance of fully internal tetrahedra, a phenomenon I refer to as lettucing. In these cases, internal layering is necessary to fill the bulk.
By contrast, when flatness is concentrated on one or two faces, for example, in a polyhedron with a large polygonal base or cap, the resulting internal tetrahedra tend to form a single radial layer, each anchored to the large flat surface. These configurations are shallow: all tetrahedra have at least one face or edge exposed.
These behaviors become increasingly common and complex as $V$ increases, due to the geometric and combinatorial expansion of allowable configurations. Thus, while heuristic rules can be proposed for internal structure at a given $V$, e.g., $4\leq V\leq 8$, such rules are often misleading, as they obscure emergent phenomena like volumetric layering that only manifest at higher vertex counts.
4.6 Example for $V=8$
$S$ (External) $T$ (Tetrahedra) $N_{i}$ (Triangles) $S_{i}$ (Segments)
0 5 4 0
0 6 6 1
0 7 8 2
0 8 10 3
1 5 4 0
1 6 6 1
1 7 8 2
2 5 4 0
2 6 6 1
3 5 4 0
4 5 4 0
5 5 4 0
6 5 4 0
From these configurations, the following heuristic rules might be proposed:
  • When $S=0$: All V4V-4 configurations are valid.
  • When $S=1$: The configuration with $T_{\max}$, $N_{i\,\max}$, and $S_{i\,\max}$ is excluded.
  • When $S=2$: The top 2 configurations are excluded.
  • When $S=k$: Configurations with $T\in[T_{\max}-(k-1),T_{\max}]$ are excluded.
  • When $S\geq V-5$: The number of valid configurations becomes narrow, and one might assume it reduces to the minimal configuration $(T_{\min},N_{i\min},S_{i\min})$. However, this assumption fails immediately with the inclusion of one additional vertex.
Consider the following two polyhedra, both with $V=9$ and $S=5$:
  • An octagon capped by eight triangles (a domed structure): flatness is concentrated on one side. It requires only $T=6$ tetrahedra, with $S_{i}=0$ and $N_{i}=5$.
  • A cube topped by a square pyramid (5 squares and 4 triangles): flatness is more evenly distributed. It requires $T=7$ tetrahedra, with $S_{i}=1$ and $N_{i}=7$.
Despite having the same surface flatness $S$, the second polyhedron requires more tetrahedra and internal gluing.
Finally, consider a polyhedron with many quadrilateral faces spread across all directions, such as the cube tunnel with $V=12$, $S=10$, and no external triangles. This structure also requires $T=T_{\min}+1$, even though surface flatness is maximal and triangle count is zero. The internal complexity in this case emerges from geometric necessity demonstrating that distributed flatness induces lettucing, even under strict constraints.
5 Inferring External Structure from Internal Decomposition and Euler’s Characteristic
This section relates the internal volumetric decomposition of a genus-0 polyhedron into tetrahedra to the combinatorics of its boundary surface. We first record an exact topological identity that holds for any tetrahedralization of a 33–ball, then specialize it to the normal–form setting used throughout this paper. We then present heuristic incidence formulas that allow one to infer $V$, $E$, and $F$ from $(T,N_{i},S,S_{i})$, and an “extended Euler-type” identity that follows under those heuristic assumptions.
5.1 Exact topological Euler link
Proposition 5.1 (Exact 3D Euler link).
Let $T$ be the number of tetrahedra in a genus-0 tetrahedralization, $N_{i}$ the number of internal gluing triangles, $E_{i}$ the number of interior edges, and $V_{i}$ the number of interior vertices. Then
$\boxed{T-N_{i}+E_{i}-V_{i}=1}$
always holds.
Proof.
Count cells in the tetrahedral complex:
  • Vertices: $V+V_{i}$ (boundary plus interior),
  • Edges: $E_{\triangle}+E_{i}$, where $E_{\triangle}$ are the edges of the triangulated boundary,
  • Faces: $N+N_{i}$, where NN is the number of boundary triangular faces (angle units),
  • 3–cells: $T$ tetrahedra.
Euler’s formula for a 33–ball gives
$(V+V_{i})-(E_{\triangle}+E_{i})+(N+N_{i})-T=1.$
Boundary Euler ($V-E_{\triangle}+F_{\triangle}=2$) with $F_{\triangle}=N$ eliminates the boundary terms, yielding
$2-E_{i}+N_{i}-T+V_{i}=1,$
which rearranges to $T-N_{i}+E_{i}-V_{i}=1$. ∎
Normal-form specialization.
In the normal form adopted here:
  • All vertices lie on the boundary $\Rightarrow V_{i}\equiv 0$,
  • Every interior edge is an internal triangulation segment $\Rightarrow E_{i}=S_{i}$.
Thus the exact identity reduces to
$\boxed{T-N_{i}+S_{i}=1}$
for all constructions considered in this manuscript.
Proposition 5.2 (Angle units and internal gluings).
Let a genus–0 polyhedron in normal form be tetrahedralized into $T$ tetrahedra with $N_{i}$ internal gluing triangles. After triangulating all boundary faces, the number of boundary triangles (angle units) is $N=F_{\triangle}=F+S=2V-4$. Counting triangle units gives
$4T-2N_{i}=N=2V-4\quad\Longrightarrow\quad\boxed{\,N_{i}=2T-V+2\,}.$
Proposition 5.3 (Exact 33D Euler link; normal-form specialization).
For any tetrahedral complex filling a 33–ball, $T-N_{i}+E_{i}-V_{i}=1$. In normal form $V_{i}=0$, and every interior edge is an interior triangulation segment; define $S_{i}:=E_{i}$. Then
$\boxed{\,T-N_{i}+S_{i}=1\,}.$
Combining with Proposition 5.2 yields the identity
$\boxed{\,T=V-3+S_{i}\,}.$
5.2 Heuristic incidence formulas for external structure
The next statement is not a general topological law; it is a combinatorial incidence calculation that assumes:
  1. 1.
    no interior Steiner vertices ($V_{i}=0$),
  2. 2.
    all internal adjacencies are along triangular faces,
  3. 3.
    $S$ counts boundary coplanarity merges,
  4. 4.
    $S_{i}$ counts internal triangulation segments completing gluing surfaces.
Heuristic Proposition 5.4 (Boundary counts from interior decomposition).
Assume a genus-zero decomposition into $T$ tetrahedra without interior Steiner vertices, where internal adjacencies occur only along triangular faces; let $N_{i}$ be the number of internal gluing triangles, $S$ the number of coplanarity merges on the boundary, and let $S_{i}$ denote the number of interior edges.
Then the boundary satisfies
$E=6T-3N_{i}-S,\qquad F=4T-2N_{i}-S,\qquad V=4T-3N_{i}+2S_{i}.$
Justification.
For a fully triangulated boundary, $F_{\triangle}=4T-2N_{i}$ and $E_{\triangle}=\frac{3}{2}F_{\triangle}=6T-3N_{i}$; $S$ coplanarity merges reduce both by $S$, giving the displayed $E$ and $F$. The $V$-formula is algebraically equivalent to the exact identities $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (Propositions 5.25.3), hence consistent. ∎
Consistency check.
Applying these expressions to two distinct polyhedra with $V=9$ and $S=5$:
Domed octagon ($T=6$, $N_{i}=5$, $S=5$, $S_{i}=0$):
$E=6\cdot 6-3\cdot 5-5=16,\quad F=4\cdot 6-2\cdot 5-5=9,\quad V=4\cdot 6-3\cdot 5+0=9.$
Cube with pyramid ($T=7$, $N_{i}=7$, $S=5$, $S_{i}=1$):
$E=6\cdot 7-3\cdot 7-5=16,\quad F=4\cdot 7-2\cdot 7-5=9,\quad V=4\cdot 7-3\cdot 7+2=9.$
In both cases: $E=16$, $F=9$, $V=9$, confirming consistency.
5.3 Heuristic extension of Euler’s characteristic
Substituting the above heuristic expressions into VE+FV-E+F gives:
$V-E+F=(4T-3N_{i}+2S_{i})-(6T-3N_{i}-S)+(4T-2N_{i}-S)=2(T-N_{i}+S_{i}).$
Heuristic 5.1 (Extended Euler-type identity).
Under the hypotheses of Proposition 5.4,
$V-E+F\;=\;2\,(T-N_{i}+S_{i}).$
Note: This is an incidence calculation that packages the boundary formulas above; it is not asserted as a general topological law. In normal form, $T-N_{i}+S_{i}=1$ (Proposition 5.3), hence VE+F=2V-E+F=2 as expected.
Reduction to the classical value in a specific family.
For the sequence $T_{k}=V-3+k$, $N_{i,k}=V-4+2k$, $S_{i,k}=k$ ($0\leq k\leq V-5$), one finds $T_{k}-N_{i,k}+S_{i,k}=1$, hence VE+F=2V-E+F=2.
6 A Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility
Section 7 formalizes a combinatorial method to assess polygon lists for enclosure feasibility, including a “flatness threshold” test. The method operates on a list of polygon types and their frequencies (e.g., 12 pentagons, 20 hexagons), without requiring coordinates or spatial modeling. It proceeds in several steps:
Step 1: Polygon Frequency:
Define the count of each polygon type $P_{k}$.
Step 2: Total Internal Angle Units (N):
Each $k$-gon contributes (k2)(k-2) angle units of 180°. Compute total:
$N=\sum_{k=3}^{n}(k-2)P_{k}$
Step 3: Vertex Count Estimate (V)
For any genus-0 polyhedron in normal form (simple 33–connected boundary; no edge bisections or interior boundary vertices), the vertex count is governed by a direct formula in terms of the face–degree counts.
Proposition 6.1.
Let $P_{k}$ denote the number of faces of degree $k$, and define
$N\;=\;\sum_{k\geq 3}(k-2)\,P_{k}.$
Then for any genus-0 polyhedron in normal form,
$\boxed{\;V\;=\;\frac{N}{2}+2\;}$
Proof.
From the face–edge incidence relation,
$2E\;=\;\sum_{k\geq 3}k\,P_{k}\;=\;\sum_{k\geq 3}(k-2)\,P_{k}\;+\;2\sum_{k\geq 3}P_{k}\;=\;N\;+\;2F.$
By Euler’s formula VE+F=2V-E+F=2, substituting $E=(N+2F)/2$ gives
$V-\frac{N+2F}{2}+F=2\quad\Rightarrow\quad V-\frac{N}{2}=2,$
and hence $V=\frac{N}{2}+2$. ∎
Corollary 6.1 (Integer feasibility check).
If NN is odd, then $V$ is non-integer and no valid genus-0 polyhedron in normal form with such a face–degree distribution can exist.
Remark (origin of the formula). An earlier derivation produced the more elaborate rational form
$V=\frac{2M_{\mathrm{total}}-N\,(2S_{\mathrm{total}}+10)}{M_{\mathrm{total}}-5N},$
with
$M=\sum_{k\geq 3}(k+2)P_{k},\quad S=\sum_{k\geq 3}(k-3)P_{k},$
arising from edge–face incidence counts and the extremal relation E=3V6SE=3V-6-S. Under the genus–0, normal–form constraints, one has M5N=4SM-5N=-4S, which causes the above expression to collapse identically to $V=\tfrac{N}{2}+2$ for all values of $S$. This shows that the vertex count is independent of flatness in this setting.
6.0.1 Derivation (for completeness)
Let $P_{k}$ be the number of kk–gonal faces, and define
$N=\sum_{k\geq 3}(k-2)P_{k},\quad M=\sum_{k\geq 3}(k+2)P_{k},\quad S=\sum_{k\geq 3}(k-3)P_{k},\quad F=\sum_{k\geq 3}P_{k}.$
From
$2E=\sum_{k\geq 3}kP_{k},\qquad 2E=3F+S,$
we find
$M=(2E)+2F,\quad N=(2E)-2F,$
and hence
$E=\frac{M+N}{4},\quad F=\frac{M-N}{4},\quad M-5N=-4(2E-3F)=-4S.$
With E=3V6SE=3V-6-S (Proposition 3.1), we have
$\frac{M+N}{4}=3V-6-S\quad\Longrightarrow\quad V=\frac{M+N+24+4S}{12}.$
Substituting M=5N4SM=5N-4S yields
$V=\frac{N}{2}+2.$
Thus the simple form follows unconditionally from the general counts.
Step 4: Triangulation Segment Count (S):
Count non-intersecting segments used in triangulation of faces:
$S_{\operatorname{total}}=\sum_{k=3}^{n}(k-3)P_{k}$
Step 5: Total Face Count (F):
$F=\sum_{k=3}^{n}P_{k}$
Step 6: Edge Count Estimate (E):
$E=3(V-2)-S_{\operatorname{total}}$
Equivalently, this identity can be rearranged to express a relationship between the total edge incidences and the number of triangulated faces: $2E+S_{\operatorname{total}}=3F$
This confirms that the triangulated surface has exactly $3F$ triangle-edge incidences, spread across edges and triangulation segments.
Step 7: Euler’s Formula Check:
VE+F=2V-E+F=2
If VE+FV-E+F does not equal to 2, then the set of polygon types does not enclose a valid 3D volume.
Step 8: Flatness Constraint:
$S_{\min}=0$
$S_{\max}=\left\{\begin{array}[]{ll}\frac{3}{2}V-\tfrac{13}{2}&\text{if }V\text{is odd}\\ &\\ \frac{3}{2}V-6&\text{if }V\text{is even}\end{array}\right$.
$S_{\operatorname{difference}}=S_{\max}-S_{\operatorname{total}}$
If $S_{\operatorname{difference}}<0$, the set of polygons is “too flat” to enclose a 3D volume.
Summary of Practical Application (Appendix D)
The combinatorial method introduced in this paper has been applied to a range of polygonal configurations, demonstrating its capacity to correctly validate known polyhedra, reject over-flattened ones, and expose structurally deceptive false positives. These worked examples include the cube, the square pyramid, and the regular dodecahedron, each of which passes all symbolic checks and matches known geometry. For example, the decomposition of the cube yields five tetrahedra which aligns with a known meshing scheme producing four right tetrahedra at its corners and a central regular tetrahedron [7]. In contrast, a configuration of 36 hexagons fails the flatness threshold, while a less intuitive case, two hexagons and four triangles, passes all algebraic constraints but appears unrealizable in $\mathbb{R}^{3}$, highlighting the symbolic method’s limits. Intriguingly, substituting two triangles with one square in that borderline case causes the structure to fail the flatness test entirely, revealing a sharp combinatorial threshold. Such substitutions may serve as diagnostic probes for detecting “brittleness” in enclosure feasibility. These examples reinforce the method’s utility as both a filtering tool for polyhedral candidates and a lens through which to explore the fine boundary between combinatorial sufficiency and geometric necessity.
7 Symmetry Between Face- and Vertex-Type Combinatorics
This section investigates the symmetry between face types and vertex valencies, drawing attention to a duality often noted but rarely operationalized in enclosure logic. In analyzing polyhedra from a combinatorial standpoint, we observe a notable parallelism between the types of faces and the types of vertices.
7.1 Face Types and Vertex Types
Each face of a polyhedron is defined by the number of its edges (or equivalently, the number of vertices around it), while each vertex is defined by the number of edges meeting at that point. These quantities are known as:
  • Face degrees: the number of edges surrounding a face,
  • Vertex valencies: the number of edges incident to a vertex.
For a polyhedron with $V$ vertices, $E$ edges, and $F$ faces, we always have the exact incidence identity
$\sum_{i=1}^{F}d_{i}\;=\;2E\;=\;\sum_{j=1}^{V}v_{j},$
where $d_{i}$ is the degree of the $i$-th face and $v_{j}$ is the valency of the $j$-th vertex. This face–vertex symmetry is emphasized in Grünbaum’s treatment of polyhedral duality [6], where vertex valencies and face degrees are shown to reflect a deeper structural reciprocity across convex forms. It reflects a basic balance: each edge is shared by two faces and connects two vertices, hence contributes twice to the totals.
7.2 Observed Symmetry and Interpretation
From empirical examination and combinatorial modeling, I notice that:
  • The range of face types (triangle, quadrilateral, pentagon, etc.) mirrors the range of vertex types (valency 3, 4, 5, etc.).
  • The minimum possible face or vertex type is generally 3 (a triangle or trivalent vertex), which is necessary for a polyhedron to be enclosed.
  • The maximum number of edges in a face or incident edges at a vertex is bounded by the number of vertices, typically less than $V$.
This observation does not suggest a strict one-to-one mapping but highlights a structural analogy:
  • When enumerating or constructing valid polyhedra, both sets of types must satisfy the same total:
    $\text{Total face degrees}=\text{Total vertex valencies}=2E.$
  • This constrains generation algorithms and supports validation strategies.
  • It provides a balanced perspective when analyzing polyhedra from the inside (vertices) or outside (faces).
Though rooted in standard topological rules, this symmetry has not been prominently emphasized in the literature as a guiding structure. It suggests an intuitive, combinatorial harmony between internal and external complexity.
7.3 Face-Type Combinatorial Constraints
The following determines valid face-type combinations in a polyhedron defined by:
  • Total number of vertices $V$,
  • External triangulation segment count $S$,
  • Total number of faces $F$,
  • Total number of edges $E$.
Face Variables and Degree Constraints.
Let the variables represent the number of faces of each type:
  • $a$: triangles (3 sides),
  • $b$: quadrilaterals (4 sides),
  • $c$: pentagons (5 sides),
  • etc., up to degree V1V-1.
Face Type Limit Rule. The face with the largest number of edges allowed in a polyhedron with $V$ vertices must have at most V1V-1 edges, since a face cannot be incident with more than V1V-1 unique vertices in a closed 3D structure.
Equations for Face-Type Combinations.
For given values of $F$, $E$, and $S$, the face-type variables must satisfy:
$a+b+c+\ldots$ $=F$ (face count)
$3a+4b+5c+\ldots$ $=2E$ (edge incidence)
$0a+1b+2c+\ldots$ $=S$ (triangulation segments).
All solutions must be non-negative integers.
Worked Example ($V=6$, $S=2$, $E=10$, $F=6$).
Allowed face types: triangles, quadrilaterals, and pentagons (up to 5 sides). Valid integer solutions:
  • $a=4$, $b=2$, $c=0$  (4 triangles, 2 quadrilaterals),
  • $a=5$, $b=0$, $c=1$  (5 triangles, 1 pentagon).
Both satisfy:
$a+b+c$ $=6$
[…]
$b+2c$ $=2.$
7.4 Vertex-Type Combinatorial Constraints
We now narrow down the valid vertex-type combinations in a polyhedron defined by:
  • $V$: total number of vertices,
  • $E$: total number of edges,
  • $S$: number of triangulation segments,
  • $T$: total number of tetrahedra.
Vertex Variables.
Let the variables represent the number of vertices of each degree type:
  • $a^{\prime}$: vertices of degree 3,
  • $b^{\prime}$: vertices of degree 4,
  • $c^{\prime}$: vertices of degree 5,
  • etc., up to degree V1V-1.
Vertex Type System of Equations.
Every valid vertex-type configuration must satisfy:
$a^{\prime}+b^{\prime}+c^{\prime}+\ldots$ $=V$ (vertex count)
$3a^{\prime}+4b^{\prime}+5c^{\prime}+\ldots$ $=2E$ (edge incidence)
$a^{\prime}+2b^{\prime}+3c^{\prime}+\ldots$ $=2E-2V$ (since $\sum_{j}(\operatorname{deg}v_{j}-2)=2E-2V$).
All variables must be non-negative integers, and the maximum degree allowed is V1V-1.
Worked Example ($V=6$, $E=10$, $S=2$, $T=3$).
Solutions include:
  • $a^{\prime}=4$, $b^{\prime}=2$, $c^{\prime}=0$,
  • $a^{\prime}=5$, $b^{\prime}=0$, $c^{\prime}=1$.
Both satisfy all constraints.
Worked Example: A Likely Non-Realizable Configuration.
Given $V=6$, $E=11$, $S=1$, $T=3$, valid solutions include:
[…]
Configs 1 and 2 appear realizable; Config 3 fails embedding. The corresponding face-type configuration is $a=6$ (triangles), $b=1$ (quadrilateral), which matches edge counts but resists geometric embedding.
7.5 Limitations and Realizability Filters
This illustrates a key limitation: satisfying combinatorial constraints does not guarantee embeddability in three-dimensional space. As Grünbaum notes, only graphs that are planar and 3-connected can be realized as convex 3D polyhedra, a result due to Steinitz [6].
To reduce false positives, one may introduce heuristic filters:
  • Bounding the number of high-valency vertices relative to the number of tetrahedra,
  • Angle feasibility checks,
  • Limits on repeated vertex incidence across tetrahedra.
Such filters are not definitive, but they help identify configurations that are combinatorially possible yet geometrically implausible.
7.6 Summary
In summary, the structural symmetry between face-type and vertex-type combinatorics provides a unifying framework. The parallel systems of equations show that both perspectives encode the same edge-incidence information while imposing complementary feasibility conditions. This dual viewpoint clarifies why some integer solutions admit realizable polyhedra while others fail, highlighting the balance between external and internal complexity in polyhedral structures.
8 On All Likely Configurations of a Polyhedron and The Limits of Combinatorial Realizability
Section 10 explores a possible route for exploring all likely configurations of a polyhedron defined by a number of vertices and derived surface flatness by utilizing  an unrestricted partition function $p(S).$ As shown in earlier sections, every valid polyhedron of $V$ vertices is characterized by two distinct combinatorial structures:
[…]
  • The internal structure: $(T,N_{i},S_{i})$. For a polyhedron of vertices V, these complements have a total number of combinations of V4V-4
I now define a function that estimates the number of valid combinations between these two sets for a given vertex count $V$. This count reflects how many internally and externally compatible sets of structures exist, without specifying the actual types of those structures (Face-types and Vertex-types).
8.1 Structure Pair Combinations
$\text{Combinations}(V)=\left\{\begin{array}[]{ll}\frac{1}{2}V^{2}-3V+5&\text{if }V\text{is even}\\ \frac{1}{2}V^{2}-3V+\tfrac{9}{2}&\text{if }V\text{is odd}\end{array}\right$.
8.2 Example Table (V = 4 to 20)
$V$ (Vertices) Estimate Total External and Internal Configuration Sets
4 1
5 2
6 5
7 8
8 13
9 18
10 25
11 32
12 41
13 50
14 61
15 72
16 85
17 98
18 113
19 128
20 145
8.2.1 Graphical Representation
Figure 4: Combinatorial sum of internal and external structure groupings as a function of vertex count.
When attempting to uncover the actual face-type and vertex-type configurations corresponding to a specific polyhedron defined by vertex count $V$, triangulation segment count $S$, and tetrahedron count $T$, the procedures described in Section 7 (Constraints on Face Type Integer Combinations and Maximum Face Degree Rule) and Section 8 (Constraints on Vertex Type Integer Combinations and Maximum Edges per Vertex Degree Rule) should be applied. These methods allow one to solve for specific polygonal face arrangements and converge to vertex degree structures that are compatible with a given global combinatorial configuration.
It is important to emphasize that satisfying all combinatorial constraints does not necessarily guarantee spatial realizability of all proposed configurations. This limitation is most apparent in the vertex-type integer combinations but may extend to other aspects of the framework as well. Some solutions that are valid in integer space may still be non-embeddable in three-dimensional geometry due to overlap or topological obstruction. Realizability remains a geometric problem. One that combinatorics alone cannot always resolve.
8.3 Estimating the Upper Bound from Flatness and Face Combinations
While previous sections emphasize compatibility between internal and external combinatorics, it is also useful to estimate the maximum number of possible polyhedral configurations based solely on the total surface angle budget. In contrast to a full combinatorial analysis of vertex and edge constraints, this method estimates the number of ways one can arrange polygonal faces on a polyhedral surface using only the flatness range of the surface. It does not take into account local realizability constraints, such as vertex incidence or internal fillability by tetrahedra. As it doesn’t account for vertex types, it serves as a partial initial upper-bound estimate of surface complexity.
8.4 Method Outline
Step 1: Compute the flatness range
Define the triangulation segment budget $S$, where $S=0$ corresponds to a surface composed entirely of triangles (i.e., zero extra flatness beyond minimal curvature), and larger values of $S$ correspond to the inclusion of increasingly flat faces (e.g., quadrilaterals, pentagons). The minimum is always $S_{\min}=0$, corresponding to the fully triangulated case. The maximum possible number of triangulation-equivalent segments $S_{\max}$ depends on the number of vertices $V$ and follows the previously mentioned empirically derived linear functions:
$S_{\max}=\left\{\begin{array}[]{ll}\frac{3}{2}V-\tfrac{13}{2}&\text{if }V\text{is odd}\\ \frac{3}{2}V-6&\text{if }V\text{is even}\end{array}\right$.
Step 2: Enumerate face-type combinations using integer partitions
Each valid configuration of polygonal face types can be expressed as an integer partition of $S$, where each polygon type contributes a fixed number of “excess segments” relative to a triangle:
  • A quadrilateral contributes 1,
  • A pentagon contributes 2,
  • And so on.
Thus, every value of $S\in[0,S_{\max}]$ corresponds to a number of integer partitions under these constraints. The total number of valid polygonal configurations at a given $V$ is then equal to the sum of these partition counts from $S_{\min}$ to $S_{\max}$. If no restriction is placed on the number of polygon sides, then this count becomes equivalent to the unrestricted partition function $p(S)$, a classical object in number theory that enumerates the number of ways to write $S$ as a sum of positive integers (ignoring order) [2]. The partition function begins:
$p(S)=\{1,1,2,3,5,7,11,15,22,30,42,\ldots\}$
Therefore, for a given vertex count $V$, the number of distinct face-type combinations is:
$\sum_{S=0}^{S_{\max}}p(S)$
For instance, for $V=5$, I compute $S_{\max}=1$, and the number of valid configurations is:
$p(0)+p(1)=1+1=2$
Step 3: Compute total number of face-type combinations
Sum all partition counts (unrestricted) across the flatness range. The resulting total gives an upper bound on the number of distinguishable surface configurations for a polyhedron with $V$ vertices, assuming only the flatness constraint.
Vertices $S_{\min}$ $S_{\max}$ Face Type Combinations (Upper Bound)
4 0 0 1
5 0 1 2
6 0 3 7
7 0 4 12
8 0 6 30
9 0 7 45
10 0 9 97
11 0 10 139
12 0 12 272
13 0 13 373
14 0 15 684
15 0 16 915
16 0 18 1597
17 0 19 2087
18 0 21 3506
19 0 22 4508
20 0 24 7338
8.5 Graphical Representation
Figure 5: Upper bound of face type combinations as a function of vertex count.
This curve provides an upper bound on the number of surface configurations derivable from face configurations for each vertex count $V$, assuming only the flatness constraint is satisfied. It does not account for limitations imposed by vertex-type configurations, or tetrahedral fillability. This is demonstrated in Appendix A, as different vertex-type configurations can exist for the same face configuration. Nonetheless, it offers a fast and interpretable way to explore surface complexity based solely on flatness and consequent face composition.
After determining the number of necessary triangles to achieve closure at a desired vertex count, each potential face-type configuration generated via this enumeration can then be filtered or corrected by the stricter combinatorial criteria, such as the one outlined in Section 7 (”A Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility”). Accordingly, adjustments, such as the addition or removal of triangular faces, can be made to align each configuration with more realizable polyhedral enclosures.
Step 4: Refine with restricted partitions
If a maximum polygon side count $k_{\max}$ is imposed (such as limiting the use of hexagons or excluding higher-order polygons) the enumeration becomes a case of restricted integer partitions, where only parts of size $\leq k_{\max}-3$ are allowed. This constraint reflects that each non-triangular face contributes a specific number of excess triangulation segments. Restricted partitions offer a more realistic estimate by excluding configurations that rely on geometrically unlikely face types. For example, limiting the count of hexagons in some configurations prevents overcounting face combinations that cannot physically tile a polyhedral surface. By summing the number of valid restricted partitions for each value of $S=0$ to $S_{\max}$, I obtain a tighter upper bound on the number of feasible face-type combinations, filtering out many false positives that are combinatorially valid but geometrically unconstructible.
9 Discussion: Internal Validity, External Limits, and Future Directions
This study has introduced a symbolic, coordinate-free framework for assessing the enclosability and internal structure of 3D polyhedra. By relying solely on integer arithmetic, triangle counts, and combinatorial rules, the method offers a new approach to reasoning about polyhedral structure without invoking geometric embedding or spatial coordinates.
While prior work has shown the computational difficulty of minimizing tetrahedra in convex polyhedra [3], the present approach, rather than aiming for minimality, characterizes the symbolic and combinatorial constraints required for enclosure and internal decomposition. A central strength of the framework presented in this study is that it leads naturally to Euler’s characteristic, one of the foundational invariants in polyhedral topology. The identity
VE+F=2V-E+F=2
emerges not by assumption, but as a direct consequence of empirically derived symbolic constraints involving internal triangulation segments, gluing surfaces, and tetrahedral decompositions. The extended form
$V-E+F=2(T-N_{i}+S_{i})$
provides a deeper lens into the interaction between internal and external structure. Based on the empirical formulas presented in this document, it was shown that for valid genus-zero decompositions,
$T-N_{i}+S_{i}=1$
which guarantees the recovery of the classical Euler result. The fact that a purely combinatorial, coordinate-free framework yields this topological invariant is not incidental; it reflects the framework’s internal consistency and structural legitimacy. Unlike abstract proofs that assume closure, this derivation is constructive: Euler’s characteristic emerges from accounting for internal tetrahedra and their gluing segments and shared triangular surfaces.
That said, while this model enforces strict symbolic coherence, it does not guarantee geometric embeddability in three-dimensional Euclidean space. Certain configurations, especially those near the maximal “flatness” threshold (i.e., where the number of triangulation segments approaches its upper bound for polyhedron of vertices $V$), may pass all combinatorial checks but still fail to realize as actual 3D polyhedra. These are symbolic false positives: structurally valid according to the framework, but geometrically unrealizable, due to overlap, crowding, or angular conflict.
This is not a weakness, but rather a natural boundary of combinatorial reasoning. It marks the limits between combinatorial sufficiency and geometric necessity. By identifying such edge cases, the framework defines the outer envelope of feasible structures and offers a direction for further refinement.
Future versions of the method may incorporate geometric heuristics, such as convexity filters, vertex angle bounds, or planarity constraints, to help exclude non-realizable cases. Computational embedding algorithms or realizability tests (based on graph-theoretic results like Steinitz’s theorem) could complement the current symbolic system, bridging toward a hybrid topological–geometric model.
Additionally, and as previously mentioned, the method is currently limited to genus-zero polyhedra, and does not yet account for higher-genus or non-manifold forms. Internal decompositions assume non-intersecting tetrahedra formed strictly through external and internal triangulation segments, without edge bisections or face subdivisions. Despite these constraints, this exploratory framework presents a symbolic heuristic for reasoning about polyhedral structure. It offers a consistent, transparent method for exploring enclosure, triangulation, and internal volume using nothing but integer logic. In doing so, it could open a pathway toward new classifications of polyhedra, and perhaps even toward insights into deeper rules that govern 3D structure.
10 Acknowledgements
The author has no formal mathematical training and, despite seeking guidance over fifteen years ago, did not receive the mentorship or institutional support needed to develop these ideas at the time. Mental health difficulties further complicated the ability to engage in collaborative research across unfamiliar disciplines, and financial constraints prevented access to supplementary education. The mathematical concepts, results, and proofs presented in this paper originate entirely from the author’s independent observations, drawings, and combinatorial reasoning, developed between his early teenage years and early twenties.
The author thanks Mr. Mohammad Chalhoub for providing programming support to produce the table of restricted integer partitions (prior to the author realizing the sequence had long been established), and Mr. Amr Tamimi for introducing him in earlier years to Excel, both contributions made many years ago. More recently, the author thanks Mr. Tiago Monteiro-Henriques and Ms. Hajar Assaad for briefly reviewing the manuscript, and Ms. Federica Manca and Ms. Anu Virtanen for assistance with illustrations.
Appendix B: Topological Comparison of Tetrahedralization Strategies and Breakdown of the Proposed Identity with Internal Steiner Points
To evaluate how different tetrahedralization strategies affect the internal structure of a cube, I analyze key topological quantities including the number of tetrahedra $T$, internal gluing triangle faces $N_{i}$, and internal triangulation segments $S_{i}$. The proposed identity
$T-N_{i}+S_{i}=1$
is observed to hold consistently across minimal and marching tetrahedra configurations, provided no internal vertices are added. However, once an internal Steiner point is introduced, particularly when it is treated as an internal vertex connected to new edges or faces, this relation fails. The following table presents a comparative breakdown of several configurations, highlighting how modeling assumptions (e.g., counting internal segments as edges or new pyramid faces as external) affect topological consistency. This analysis reinforces the significance of preserving external-only structure in minimal decompositions and offers a reference for evaluating topological soundness in more complex triangulations.
Configuration $V$ $E$ $F$ $T$ $N_{i}$ $S_{i}$ EFE{-}F $2T{-}N_{i}$ VE+FV{-}E{+}F $T{-}N_{i}{+}S_{i}$
Minimal 8 12 6 5 4 0 6 6 2 1
Marching Tetrahedralization 8 12 6 6 6 1 6 6 2 1
Steiner not counted as vertex 8 12 6 12 18 0 6 6 2 -6
Steiner counted as vertex 9 12 6 12 18 0 6 6 3 -6
Steiner counted + edges 9 20 6 12 18 0 14 6 -5 -6
Steiner counted + edges + faces 9 20 18 12 6 0 2 18 7 6
This comparison demonstrates that the proposed identity is robust when internal structure arises solely from external constraints. In the minimal 5-tetrahedra decomposition, the identity holds exactly with no added internal triangulation segments. In the marching tetrahedralization strategy, although the identity still holds, the cube is bisected into two hexahedra by introducing an internal quadrilateral face. This increases internal complexity: one additional tetrahedron is required ($T$ increases by 1), along with two new internal gluing triangles ($N_{i}$) and one internal triangulation segment ($S_{i}$). Nonetheless, the identity is preserved due to the balance between these added quantities. However, once a Steiner point is introduced inside the volume and treated as a vertex connected by new internal edges or faces, the identity fails. This breakdown reflects a fundamental shift in the topological model: the internal structure is no longer a consequence of surface triangulation alone, but instead includes additional volumetric elements. These results emphasize the importance of distinguishing between surface-conforming decompositions and those requiring internal augmentation when analyzing or designing tetrahedral meshes.
Importantly, the breakdown observed in the proposed identity when internal Steiner points are introduced is not isolated. The classical Euler characteristic
$\chi=V-E+F=2$
for convex polyhedra also fails to hold in such configurations. For instance, in the case where internal segments and internal pyramid faces are explicitly counted (“Steiner counted + edges + faces”), I obtain
$\chi=9-20+18=7.$
This discrepancy highlights that the underlying structure is no longer homeomorphic to a topological sphere. Once internal vertices, edges, or faces are introduced and treated as part of the overall complex, both Euler’s characteristic and the proposed identity cease to apply in their original forms. This caveat highlights the importance of distinguishing between boundary-based and volumetrically augmented models when performing topological accounting. Now, compare with Appendix C for a case where invariance is preserved under surface-Steiner points.
Appendix C: Invariance of the Proposed Identity for Polyhedra with Surface-Embedded Edge Steiner Points
To illustrate the invariance of the identity $T-N_{i}+S_{i}=1$ under different tetrahedral decompositions, including those involving Steiner points, consider a single tetrahedron with one Steiner point placed on each of three edges, each edge chosen such that the three Steiner points lie  on the same face.
Decomposition A (Plane Cut Through Three Steiner Points):
Consider the plane passing through the three Steiner points. This plane divides the tetrahedron into two polyhedral regions:
  • One region is a tetrahedron.
  • The other is a five-faced polyhedron bounded by three quadrilateral faces and two triangular faces, with six vertices in total.
This five-faced polyhedron can be further decomposed into three tetrahedra, yielding a total of four tetrahedra. The tetrahedra are glued pairwise along three internal triangular faces, and the only Steiner points used are the original three placed on the edges.
Thus, the identity is satisfied:
$T=4,N_{i}=3,S_{i}=0\Rightarrow T-N_{i}+S_{i}=4-3+0=1$
Decomposition B (Edge-Based Slicing Through Steiner Points):
Alternatively, one can sequentially slice the tetrahedron through one Steiner point at a time, joining each to the two original vertices that define its edge, and to one of the remaining original vertices, forming tetrahedra that share triangular faces. This approach also produces four tetrahedra, glued across the same number of internal triangular interfaces (three), and using the same three Steiner points.
Again, the identity holds:
$T-N_{i}+S_{i}=4-3+0=1$
This example shows that despite the different topology of the intermediate polyhedral shapes, the identity is preserved, emphasizing its combinatorial invariance across Steiner-supported decompositions.
Appendix D: Worked Examples Demonstrating the Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility Presented in Section 7
This appendix presents worked examples that demonstrate the combinatorial method. It includes three examples that are combinatorially and geometrically valid, one example that is too flat to enclose, and an example of a likely false positive, a case where the set of polygons is determined symbolically valid but likely unrealizable.
Worked Example: Four Triangles and One Quadrilateral (Square Pyramid)
  • Step 1: Polygons: 1 square ($P_{4}=1$), 4 triangles ($P_{3}=4$)
  • Step 2: $N=6$
  • Step 3: $V=\frac{6}{2}+2=5$
  • Step 4: $S=1$
  • Step 5: $F=5$
  • Step 6: $E=3\cdot(5-2)-1=8$
  • Step 7: Euler check: 58+5=25-8+5=2
  • Step 8: $S_{\max}=\frac{3}{2}\cdot 5-\tfrac{13}{2}=2$; $S_{\mathrm{difference}}=1$ (passes flatness test)
Worked Example: Six Squares (The Cube)
  • Step 1: Polygons: 6 squares or quadrilaterals ($P_{4}=6$)
  • Step 2: $N=(4-2)\cdot 6=12$
  • Step 3: $V=\frac{12}{2}+2=8$
  • Step 4: $S=(4-3)\cdot 6=6$
  • Step 5: $F=6$
  • Step 6: $E=3\cdot(8-2)-6=12$
  • Step 7: Euler check: 812+6=28-12+6=2
  • Step 8: $S_{\max}=\frac{3}{2}\cdot 8-6=6$; $S_{\mathrm{difference}}=0$ (passes flatness test)
Worked Example: 12 Pentagons (The Dodecahedron)
  • Step 1: Polygons: 12 pentagons ($P_{5}=12$)
  • Step 2: $N=(5-2)\cdot 12=36$
  • Step 3: $V=\frac{36}{2}+2=20$
  • Step 4: $S=(5-3)\cdot 12=24$
  • Step 5: $F=12$
  • Step 6: $E=3\cdot(20-2)-24=30$
  • Step 7: Euler check: 2030+12=220-30+12=2
  • Step 8: $S_{\max}=\frac{3}{2}\cdot 20-6=24$; $S_{\mathrm{difference}}=0$ (passes flatness test)
Worked Example: 36 Hexagons
  • Step 1: Polygons: 36 hexagons ($P_{6}=36$)
  • Step 2: $N=144$
  • Step 3: $V=\frac{144}{2}+2=74$
  • Step 4: $S=108$
  • Step 5: $F=36$
  • Step 6: $E=3\cdot(74-2)-108=108$
  • Step 7: Euler check: 74108+36=274-108+36=2
  • Step 8: $S_{\max}=\frac{3}{2}\cdot 74-6=105$; $S_{\mathrm{difference}}=-3$ (fails flatness test)
  • Conclusion: A set of 36 hexagons cannot produce an enclosed polyhedron.
Worked Example: Two Hexagons and Four Triangles (False Positive)
  • Step 1: Polygons: 2 hexagons ($P_{6}=2$), 4 triangles ($P_{3}=4$)
  • Step 2: $N=(6-2)\cdot 2+(3-2)\cdot 4=8+4=12$
  • Step 3: $V=\frac{12}{2}+2=8$
  • Step 4: $S=(6-3)\cdot 2+(3-3)\cdot 4=6+0=6$
  • Step 5: $F=6$
  • Step 6: $E=3\cdot(8-2)-6=12$
  • Step 7: Euler check: VE+F=812+6=2V-E+F=8-12+6=2 (passes)
  • Step 8: $S_{\max}=\frac{3}{2}\cdot 8-6=6$, $S_{\mathrm{difference}}=0$ (passes flatness)
  • Conclusion: This configuration passes all combinatorial checks: Euler’s formula holds, edge and face counts are consistent, the triangulation segment count meets the flatness threshold, and a minimal internal tetrahedral decomposition is arithmetically feasible. However, no known convex or non-convex polyhedron with exactly two hexagons and four triangles exists in the literature. The structure cannot be embedded in $\mathbb{R}^{3}$ as a genus-zero surface with non-self-intersecting faces. This mismatch illustrates the boundary between symbolic sufficiency and geometric feasibility. While the framework successfully captures algebraic consistency, it cannot, by design, account for spatial constraints such as embeddability or convex closure. This example thus serves as a clear case where realizability remains unresolved. No known convex or non-convex realization exists.
    When substituting two of the triangles with one quadrilateral, keeping the total face count at six, the resulting configuration becomes too flat to enclose. This failure highlights how even minimal changes in polygonal composition can push a structure past the limits of combinatorial enclosure. Such sensitivity underscores the brittleness of symbolic tetrahedral enclosure near flatness thresholds. The substitution of two triangles with a quadrilateral may therefore serve as a practical diagnostic for testing the combinatorial resilience of enclosure candidates. Identifying and interpreting such tipping points offers a pathway for refining this framework and integrating embeddability heuristics in future work.
Appendix E: Combinatorial Identities and Bounds
This appendix records (i) proved external and topological identities used in the main text, and (ii) the empirical/heuristic identities and bounds collected from worked decompositions. All variables are symbolic and coordinate-free. Genus 0 and normal-form assumptions apply where stated. All linear “max” values quoted below are valid within the SALT+MIE class (Definition 2.1). Non–SALT triangulations, such as dense pulling triangulations of cyclic polytopes, can exceed these bounds by creating quadratically many tetrahedra without adding interior vertices.
E.1 Proved External Identities and Bounds
Exact external formulas for genus-0 polyhedra in normal form:
$E$ $=3V-6-S$ (E.1)
$F$ $=2V-4-S$ (E.2)
where
$S\;=\;\sum_{f}(\operatorname{deg}f-3)\quad\text{(flatness, total deviation from triangulation).}$
From the vertex degree constraint $\operatorname{deg}(v)\geq 3$, one obtains the sharp bounds:
Even $V$:
$0$ $\leq S\leq\frac{3V}{2}-6,$ (E.3a)
$\frac{3V}{2}$ $\leq E\leq 3V-6,$ (E.4aE)
$\frac{V}{2}+2$ $\leq F\leq 2V-4.$ (E.4aF)
Odd $V$:
$0$ $\leq S\leq\tfrac{3V}{2}-\tfrac{13}{2}$ (E.3b)
$\tfrac{3V+1}{2}$ $\leq E\leq 3V-6$ (E.4bE)
$\tfrac{V}{2}+\tfrac{5}{2}$ $\leq F\leq 2V-4$ (E.4bF)
Extremal Values Summary (genus-0, normal form)
$V$ parity $S_{\max}$ $E_{\min}$ $E_{\max}$ $F_{\min}$ $F_{\max}$
Even $V$ $\dfrac{3V}{2}-6$ $\dfrac{3V}{2}$ 3V63V-6 $\dfrac{V}{2}+2$ 2V42V-4
Odd $V$ $\dfrac{3V}{2}-\tfrac{13}{2}$ $\dfrac{3V+1}{2}$ 3V63V-6 $\dfrac{V}{2}+\tfrac{5}{2}$ 2V42V-4
Angle-unit bookkeeping (fully triangulated boundary):
$F_{\triangle}$ $=F+S\;=\;2V-4$ (E.5)
$F_{\triangle}$ $=4T-2N_{i}$ (E.6)
(Equation (E.6) counts triangle units: each tetrahedron contributes 44, each internal gluing triangle cancels 22.)
E.2 Proved Topological Identity for 3D Complexes
For any tetrahedralization of a 33–ball (genus–0 volume) with $V_{i}$ interior vertices and $E_{i}$ interior edges,
$T-N_{i}+E_{i}-V_{i}$ $=1$ (E.7)
(Prop. 5.1). In particular, when $V_{i}=0$ (normal form, no interior vertices),
$T-N_{i}+E_{i}$ $=1.$ (E.8)
E.3 Bridge to the Heuristic Extended Euler Form
A heuristic identity used in the text is
$V-E+F$ $=2\,(T-N_{i}+S_{i})$ (E.9h)
and the SALT ladder case asserts
$T-N_{i}+S_{i}$ $=1.$ (E.10h)
These coincide with (E.7)–(E.8) whenever $V_{i}=0$ and $S_{i}$ agrees with the interior–edge count $E_{i}$ (as is the case in SALT+MIE constructions). Thus $T-N_{i}+S_{i}=T-N_{i}+E_{i}=1$ in that restricted setting.
Core Equations (proved unless marked heuristic)
(E.1) $V-E+F=2\quad\text{(Euler, genus $0$)}$ (1)
(E.9h) $V-E+F=2(T-N_{i}+S_{i})\quad\text{(heuristic)}$ (2)
(E.10h) $T-N_{i}+S_{i}=1\quad\text{(SALT ladder case)}$ (3)
(E.11r) $N_{i}=2T-V+2\quad\text{(exact in normal form (no interior vertices)}$ (4)
[…]
Restricted SALT+MIE ranges (heuristic / construction–based)
(E.13r) $T_{\min}=V-3$ (6)
(E.14r) $T_{\max}=2(V-4)$ (7)
(E.15r) $N_{i,\min}=V-4$ (8)
(E.16r) $N_{i,\max}=3(V-4)-2$ (9)
Derived Relationships (algebraic within SALT+MIE)
The following are algebraic consequences of (E.13r)–(E.16r); they are not independent results.
(E.17r) $T_{\max}=2N_{i,\min}$ (10)
(E.18r) $N_{i,\max}=\tfrac{3}{2}T_{\max}-2$ (11)
(E.19r) $N_{i,\max}=2N_{i,\min}+(V-6)$ (12)
(E.20r) $N_{i,\max}-T_{\max}=N_{i,\min}-2$ (13)
(E.21r) $N_{i,\max}-T_{\max}=T_{\min}-3$ (14)
(E.22r) $N_{i,\min}=T_{\min}-1$ (15)
(E.23r) $2N_{i,\max}-2T_{\max}=N_{i,\min}+T_{\min}-5$ (16)
Scope note. Equations (E.1)–(E.8) are exact and universal topological identities. Equations (E.9h)–(E.23r) hold only heuristically or within the restricted SALT+MIE construction class.